KyledaGreat:

In chemistry, the pH of a solution is a measure of the acidity or alkalinity of a solution. Water has a pH of 7 and, in general, acids have a pH less than 7 and alkaline solutions have a pH greater than 7. Find the pH of a solution with a hydronium ion concentration of 5.1×10^−6 moles/liter. Round your answer to two decimal places, if necessary. pH =

KyledaGreat:

@vocaloid

Vocaloid:

-log(hydronium concentration) = -log(5.1×10^−6), plug and round

KyledaGreat:

4.24542105

Vocaloid:

I got something a little different when I plugged it into my calculator: 5.29?

KyledaGreat:

you're correct

KyledaGreat:

Consider the following functions. f(x)=|x| and g(x)= −4x Find the formula for (f/g)(x) and simplify your answer. (f/g) (x) =

Vocaloid:

f/g means divide f(x) / g(x)

KyledaGreat:

oh ok

KyledaGreat:

f(x) / g(x) = -4x right ?

Vocaloid:

f(x)=|x| and g(x)= −4x f(x) = |x| / (-4x) you can't really remove the absolute value sign without changing the function, so I would just leave it like that if you can

KyledaGreat:

oh ok i thought you would

KyledaGreat:

0

Vocaloid:

wait where is 0 coming from just divide the two functions to get |x| / (-4x), that's it, it's only asking for f/g

KyledaGreat:

i'm not sure how to divide the two functions together

Vocaloid:

|x| / (-4x) is as simple as you can get it, try that first

KyledaGreat:

-|x|/4x

Vocaloid:

yeah try that and see if they'll take that

KyledaGreat:

no

KyledaGreat:

it said The answer you have submitted, −x/4x, is incorrect. It appears you have made an error in one or more of the following areas: Dividing f(x) by g(x). Combining like terms, if necessary. You may want to check your work. For a step-by-step guide to solving this problem, select Step By Step.

Vocaloid:

... strange you can't really divide |x| by x because they're not the same thing. can you take a screencap of the problem?

KyledaGreat:

1 attachment
Vocaloid:

f(x) has an absolute value sign. can you enter |x| in the numerator instead?

Vocaloid:

the || are crucial, |x| is not the same as x

KyledaGreat:

i can't put || in

Vocaloid:

control shift \ didn't work? https://www.hawkeslearning.com/documents/KeyboardShortcuts.pdf

KyledaGreat:

no , i'll show you the answer

KyledaGreat:

1 attachment
KyledaGreat:

it doesn't make sense that i couldn't use || tho for the answer

KyledaGreat:

it doesn't make sense that i couldn't use || tho for the answer

KyledaGreat:

Consider the following functions. f(x)=√x+1 and g(x) = x2^/3 Find the formula for (fg)(x) and simplify your answer. (f/g)(x) =

Vocaloid:

it's similar to the other one, you take f(x) and divide it by g(x) can you try taking a screengrab of this one too?

KyledaGreat:

1 attachment
Vocaloid:

how many tries do you get? because for this one, you can't really simplify beyond f(x) / g(x) = sqrt(x+1) / x^(2/3)

KyledaGreat:

1 time. i tried to do a similar try but i could skip it

Vocaloid:

I don't think you can simplify beyond \[\frac{ \sqrt{x+1} }{ x^{2/3} }\]but we can always skip and go back to it I guess

KyledaGreat:

it's okay , i already skipped it now

KyledaGreat:

it's complicated